# Inversion of adjunction for quotient singularities III: semi-invariant case

### 10 Dec 2023 v1

@@ 4. Order calculations
$\left(\bigwedge^{c}\left(I_{2}/I_{2}^{2}\right)\otimes_{S_{2}}S_{4}\right)\otimes_{S_{4}}\Omega_{S_{4}/k[t^{1/d}]}^{N-c}\to\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}\otimes_{R[t^{1/d}]}S_{4}.$
The $S_{4}$-module $\left(\bigwedge^{c}\left(I_{2}/I_{2}^{2}\right)\otimes_{S_{2}}S_{4}\right)\otimes_{S_{4}}\Omega_{S_{4}/k[t^{1/d}]}^{N-c}$ is generated by the elements of the form
@@
±$\pm t^{w_{\gamma}({\bf f})}\cdot\Delta\cdot dx_{1}\wedge\cdots\wedge dx_{N},$
@@
$t^{w_{\gamma}({\bf f})}\cdot\operatorname{Fitt}^{n}\left(\Omega_{B_{4}/k[t^{1/d}]}\right)\left(\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}\otimes_{R[t^{1/d}]}S_{4}\right),$
@@
By composing $\Omega^{n}_{B_{1}/k[t^{1/d}]}\big{|}_{B_{4}}\to\Omega^{n}_{B_{4}/k[t^{1/d}]}$ and $\Omega^{n}_{B_{4}/k[t^{1/d}]}\to L_{4}$ in Lemma 4.3(1), we obtain a natural morphism $\Omega^{n}_{B_{1}/k[t^{1/d}]}\big{|}_{B_{4}}\to L_{4}$. Then since $L_{4}$ is generated by one global section, we define an ideal sheaf $\mathfrak{n}^{\prime}_{1,4}\subset\mathcal{O}_{B_{4}}$ by
$\operatorname{Im}\left(\Omega^{n}_{B_{1}/k[t^{1/d}]}\big{|}_{B_{4}}\to L_{4}\right)=\mathfrak{n}^{\prime}_{1,4}\otimes_{\mathcal{O}_{B_{4}}}L_{4}.$
@@
We have the following equality of ideal sheaves on $B_{4}$:
$t^{r\cdot\operatorname{age}(\gamma)}\left(\mathfrak{n}_{r,B}\mathcal{O}_{B_{4}}\right)=t^{r\cdot w_{\gamma}({\bf f})}\left(\mathfrak{n}^{\prime}_{1,4}\right)^{r}.$
@@
By the definition of $\mathfrak{n}^{\prime}_{1,4}$, we have
$\operatorname{Im}\left(t^{r\cdot w_{\gamma}({\bf f})}\left(\Omega^{n}_{B_{1}/k[t^{1/d}]}\right)^{\otimes r}\big{|}_{B_{4}}\to L_{4}^{\otimes r}\right)=t^{r\cdot w_{\gamma}({\bf f})}\left(\mathfrak{n}^{\prime}_{1,4}\right)^{r}\otimes L_{4}^{\otimes r}.$
@@
$\operatorname{Im}\left(t^{r\cdot w_{\gamma}({\bf f})}\left(\Omega^{n}_{B_{1}/k[t^{1/d}]}\right)^{\otimes r}\big{|}_{B_{4}}\to\omega_{B_{1}/k[t^{1/d}]}^{[r]}\big{|}_{B_{4}}\right)=t^{r\cdot w_{\gamma}({\bf f})}\left(\mathfrak{n}_{r,B}\mathcal{O}_{B_{4}}\right)\otimes\omega_{B_{1}/k[t^{1/d}]}^{[r]}\big{|}_{B_{4}}.$
@@
$\operatorname{Im}\left(t^{r\cdot w_{\gamma}({\bf f})}\left(\omega_{B_{1}/k[t^{1/d}]}^{[r]}\big{|}_{B_{4}}\right)\to L_{4}^{\otimes r}\right)=t^{r\cdot\operatorname{age}(\gamma)}L_{4}^{\otimes r}.$
@@
$\operatorname{Im}\left(t^{r\cdot w_{\gamma}({\bf f})}\left(\Omega^{n}_{B_{1}/k[t^{1/d}]}\right)^{\otimes r}\big{|}_{B_{4}}\to L_{4}^{\otimes r}\right)=t^{r\cdot\operatorname{age}(\gamma)}\left(\mathfrak{n}_{r,B}\mathcal{O}_{B_{4}}\right)\otimes L_{4}^{\otimes r},$
@@
Note that $\overline{B}^{(\gamma)}\times_{\operatorname{Spec}k[t]}\operatorname{Spec}k[t^{1/d}]=B_{4}$. We pick a lift $\beta:\operatorname{Spec}k[[t^{1/d}]]\to B_{4}$ of $\alpha:\operatorname{Spec}k[[t]]\to\overline{B}^{(\gamma)}$. For an ideal sheaf $\mathfrak{a}\subset\mathcal{O}_{B_{4}}$, we define $\operatorname{ord}_{\beta}(\mathfrak{a})$ as follows: For the corresponding ring homomorphism $\beta^{*}:\mathcal{O}_{B_{4}}\to k[[t^{1/d}]]$, we have $\beta^{*}(\mathfrak{a})k[[t^{1/d}]]=\bigl{(}t^{\frac{\ell}{d}}\bigr{)}$ for some $\ell\in\mathbb{Z}_{\geq 0}$. Then, we define $\operatorname{ord}_{\beta}(\mathfrak{a}):=\frac{\ell}{d}$.
@@
Since $\operatorname{Fitt}^{n}\left(\Omega_{B_{4}/k[t^{1/d}]}\right)=\operatorname{Jac}_{\overline{B}^{(\gamma)}/k[t]}\mathcal{O}_{B_{4}}$, we have
@@
$\operatorname{ord}_{\beta}\left(\operatorname{Fitt}^{n}\left(\Omega_{B_{4}/k[t^{1/d}]}\right)\right)=\operatorname{ord}_{\alpha}\left(\operatorname{Jac}_{\overline{B}^{(\gamma)}/k[t]}\right)<\infty.$
Thus by the same argument as in [NS22]*Lemma 4.6, we have
$\frac{a}{d}+\operatorname{ord}_{\beta}\left(\operatorname{Fitt}^{n}\left(\Omega_{B_{4}/k[t^{1/d}]}\right)\right)=\operatorname{ord}_{\beta}(\mathfrak{n}^{\prime}_{1,4}),$
where $a$ is the length of
@@
$\operatorname{Coker}\left(\beta^{*}\left(\Omega_{B_{1}/k[t^{1/d}]}\big{|}_{B_{4}}\right)\to\left(\beta^{*}\Omega_{B_{4}/k[t^{1/d}]}\right)\big{/}\left(\beta^{*}\Omega_{B_{4}/k[t^{1/d}]}\right)_{\rm tor}\right)$
as a $k[t^{1/d}]$-module. Since
@@
$\operatorname{Coker}\left(\beta^{*}\left(\Omega_{B_{1}/k[t^{1/d}]}\big{|}_{B_{4}}\right)\to\left(\beta^{*}\Omega_{B_{4}/k[t^{1/d}]}\right)\big{/}\left(\beta^{*}\Omega_{B_{4}/k[t^{1/d}]}\right)_{\rm tor}\right)$
@@
we have $\frac{a}{d}=\operatorname{ord}_{\alpha}\left(\operatorname{jac}_{\overline{\mu}_{\gamma}}\right)$ Therefore, the assertion follows from Lemma 4.5. ∎
We define a $k[t]$-scheme $B_{5}$ by
$B_{5}$ $:=\operatorname{Spec}\left(R^{C_{\gamma}}[t^{1/d}]\big{/}\left(I_{\overline{B}^{(\gamma)}}\cap R^{C_{\gamma}}[t]\right)R^{C_{\gamma}}[t^{1/d}]\right)$
@@
Then, we have a natural morphism $B_{4}\to B_{5}$.
@@
By composing $\Omega^{n}_{B_{5}/k[t^{1/d}]}\big{|}_{B_{4}}\to\Omega^{n}_{B_{4}/k[t^{1/d}]}$ and $\Omega^{n}_{B_{4}/k[t^{1/d}]}\to L_{4}$ in Lemma 4.3(1), we obtain a natural morphism $\Omega^{n}_{B_{5}/k[t^{1/d}]}\big{|}_{B_{4}}\to L_{4}$. Then, we denote by $\mathfrak{n}^{\prime}_{5,4}$ the ideal sheaf on $B_{4}$ satisfying
$\operatorname{Im}\left(\Omega^{n}_{B_{5}/k[t^{1/d}]}\big{|}_{B_{4}}\to L_{4}\right)=\mathfrak{n}^{\prime}_{5,4}\otimes_{\mathcal{O}_{B_{4}}}L_{4}.$
@@
There exists a $C_{\gamma}$-invariant ideal sheaf $\mathfrak{n}^{\prime}_{p}$ on $\overline{B}^{(\gamma)}$ such that $\mathfrak{n}^{\prime}_{5,4}=\mathfrak{n}^{\prime}_{p}\mathcal{O}_{B_{4}}$.
@@
Take elements $g_{1},\ldots,g_{\ell}\in R^{C_{\gamma}}$ satisfying $R^{C_{\gamma}}=k[g_{1},\ldots,g_{\ell}]$. For a subset $J\subset\{1,\ldots,\ell\}$ with #$\#J=N-c$, we define $\Delta_{J}\in R[t]$ as the determinant of the Jacobian matrix (of size $N$) with respect to $\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}$ for $1\leq i\leq c$ and $g_{i}$ for $i\in J$, and $\partial x_{j}$ for $1\leq j\leq N$. Note that $\Delta_{J}$ is defined up to sign. We define an ideal $\mathfrak{n}^{\prime}_{p}\subset R[t]$ as
@@
$\mathfrak{n}^{\prime}_{p}:=\bigl{(}\Delta_{J}\ \big{|}\ \text{$J\subset\{1,\ldots,\ell\}$ with $\#J=N-c$}\bigr{)}.$
For any $\alpha\in C_{\gamma}$, by Lemma 3.4(1), we have
$\alpha\left(\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\right)=\xi^{dw_{\alpha}(f_{i})}\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}.$
@@
Furthermore, note that $\alpha$ maps $x_{i}$’s to their $k$-linear independent $k$-linear combinations. Therefore, for each $J$, we have
$\xi^{dw_{\alpha}({\bf f})}\cdot\Delta_{J}\cdot dx_{1}\wedge\cdots\wedge dx_{N}$
@@
$=\xi^{dw_{\alpha}({\bf f})}\cdot\bigl{(}\wedge_{i\in J}\ dg_{i}\bigr{)}\wedge\left(\wedge_{1\leq i\leq c}\ d\left(\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\right)\right)$
@@
$=\bigl{(}\wedge_{i\in J}\ d(\alpha(g_{i}))\bigr{)}\wedge\left(\wedge_{1\leq i\leq c}\ d\left(\alpha\left(\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\right)\right)\right)$
@@
$=\alpha\left(\bigl{(}\wedge_{i\in J}\ dg_{i}\bigr{)}\wedge\left(\wedge_{1\leq i\leq c}\ d\left(\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\right)\right)\right)$
$=\alpha\left(\Delta_{J}\cdot dx_{1}\wedge\cdots\wedge dx_{N}\right)$
$=\alpha\left(\Delta_{J}\right)\cdot d(\alpha(x_{1}))\wedge\cdots\wedge d(\alpha(x_{N}))$
$=\alpha\left(\Delta_{J}\right)\cdot s\cdot dx_{1}\wedge\cdots\wedge dx_{N}$
@@
for some $s\in k^{\times}$. Hence, we have $\alpha(\Delta_{J})=s^{\prime}\cdot\Delta_{J}$ for some $s^{\prime}\in k$. Hence, we conclude that $\mathfrak{n}^{\prime}_{p}$ is $C_{\gamma}$-invariant.
@@
We shall prove that this $\mathfrak{n}^{\prime}_{p}$ satisfies the assertion $\mathfrak{n}^{\prime}_{5,4}=\mathfrak{n}^{\prime}_{p}\mathcal{O}_{B_{4}}$. We set
$S_{5}:=R^{C_{\gamma}}[t^{1/d}]\big{/}\left(I_{\overline{B}^{(\gamma)}}\cap R^{C_{\gamma}}[t]\right)R^{C_{\gamma}}[t^{1/d}].$
@@
We also use the notation in the proof of Lemma 4.3. By the proof of Lemma 4.3(1), the considered map $\Omega^{n}_{B_{5}/k[t^{1/d}]}\big{|}_{B_{4}}\to L_{4}$ is corresponding to

### 30 Apr 2026 v2

@@ 4. Order calculations
(iii) $\left(\bigwedge^{c}\left(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\right)\otimes_{S_{\overline{B}_{k[t^{1/d}]}}}S_{\overline{B}^{(\gamma)}_{1/d}}\right)$ $\otimes_{S_{\overline{B}^{(\gamma)}_{1/d}}}\Omega_{S_{\overline{B}^{(\gamma)}_{1/d}}/k[t^{1/d}]}^{N-c}$
$\longrightarrow$ $\ \Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}\otimes_{R[t^{1/d}]}S_{\overline{B}^{(\gamma)}_{1/d}}$.
The $S_{\overline{B}^{(\gamma)}_{1/d}}-$module
$\left(\bigwedge^{c}\left(I_{\overline{B}_{k[t^{1/d}]}}/I^{2}_{\overline{B}_{k[t^{1/d}]}}\right)\otimes_{S_{\overline{B}_{k[t^{1/d}]}}}S_{\overline{B}^{(\gamma)}_{1/d}}\right)\otimes_{S_{\overline{B}^{(\gamma)}_{1/d}}}\Omega_{S_{\overline{B}^{(\gamma)}_{1/d}}/k[t^{1/d}]}^{N-c}$
is generated by the elements of the form
@@
$\pm t^{w_{\gamma}({\bf f})}\cdot\Delta\cdot dx_{1}\wedge\cdots\wedge dx_{N},$
@@
$t^{w_{\gamma}({\bf f})}\cdot\operatorname{Fitt}^{n}\Bigl(\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)\Bigl(\Omega_{R[t^{1/d}]/k[t^{1/d}]}^{N}\otimes_{R[t^{1/d}]}S_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)$,
@@
By composing $\Omega^{n}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}$ and $\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\to L_{\overline{B}^{(\gamma)}_{1/d}}$ in Lemma 4.3(1), we obtain a natural morphism $\Omega^{n}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to L_{\overline{B}^{(\gamma)}_{1/d}}$. Then since $L_{\overline{B}^{(\gamma)}_{1/d}}$ is generated by one global section, we define an ideal sheaf $\mathfrak{n}^{\prime}_{\overline{B}^{(\gamma)}_{1/d}}\subset\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}$ by
$\operatorname{Im}\left(\Omega^{n}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to L_{\overline{B}^{(\gamma)}_{1/d}}\right)=\mathfrak{n}^{\prime}_{\overline{B}^{(\gamma)}_{1/d}}\otimes_{\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}}L_{\overline{B}^{(\gamma)}_{1/d}}$.
@@
We have the following equality of ideal sheaves on $\overline{B}^{(\gamma)}_{1/d}$:
$t^{r\cdot\operatorname{age}(\gamma)}\Bigl(\mathfrak{n}_{r,B}\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)=t^{r\cdot w_{\gamma}({\bf f})}\Bigl(\mathfrak{n}^{\prime}_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)^{r}.$
@@
By the definition of $\mathfrak{n}^{\prime}_{\overline{B}^{(\gamma)}_{1/d}}$, we have
$\operatorname{Im}\left(t^{r\cdot w_{\gamma}({\bf f})}\Bigl(\Omega^{n}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\Bigr)^{\otimes r}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to L_{\overline{B}^{(\gamma)}_{1/d}}^{\otimes r}\right)=t^{r\cdot w_{\gamma}({\bf f})}\Bigl(\mathfrak{n}^{\prime}_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)^{r}\otimes L_{\overline{B}^{(\gamma)}_{1/d}}^{\otimes r}.$
@@
$\operatorname{Im}$ $\left(t^{r\cdot w_{\gamma}({\bf f})}\Bigl(\Omega^{n}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\Bigr)^{\otimes r}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to\omega_{B_{k[t^{1/d}]}/k[t^{1/d}]}^{[r]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\right)$
$=t^{r\cdot w_{\gamma}({\bf f})}\Bigl(\mathfrak{n}_{r,B}\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\otimes\omega_{B_{k[t^{1/d}]}/k[t^{1/d}]}^{[r]}\big|_{\overline{B}^{(\gamma)}_{1/d}}$.
@@
$\operatorname{Im}\left(t^{r\cdot w_{\gamma}({\bf f})}\Bigl(\omega_{B_{k[t^{1/d}]}/k[t^{1/d}]}^{[r]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\to L_{\overline{B}^{(\gamma)}_{1/d}}^{\otimes r}\right)=t^{r\cdot\operatorname{age}(\gamma)}L_{\overline{B}^{(\gamma)}_{1/d}}^{\otimes r}.$
@@
$\operatorname{Im}$ $\left(t^{r\cdot w_{\gamma}({\bf f})}\Bigl(\Omega^{n}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\Bigr)^{\otimes r}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to L_{\overline{B}^{(\gamma)}_{1/d}}^{\otimes r}\right)$
$=t^{r\cdot\operatorname{age}(\gamma)}\Bigl(\mathfrak{n}_{r,B}\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\otimes L_{\overline{B}^{(\gamma)}_{1/d}}^{\otimes r},$
@@
Note that $\overline{B}^{(\gamma)}\times_{\operatorname{Spec}k[t]}\operatorname{Spec}k[t^{1/d}]=\overline{B}^{(\gamma)}_{1/d}$. Let $\beta:\operatorname{Spec}k[[t^{1/d}]]\rightarrow\overline{B}_{1/d}^{(\gamma)}$ be the unique lift of $\alpha:\operatorname{Spec}k[[t]]\rightarrow\overline{B}^{(\gamma)}$ induced by the base change. For an ideal sheaf $\mathfrak{a}\subset\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}$, we define $\operatorname{ord}_{\beta}(\mathfrak{a})$ as follows: For the corresponding ring homomorphism $\beta^{*}:\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}\to k[[t^{1/d}]]$, we have $\beta^{*}(\mathfrak{a})k[[t^{1/d}]]=\bigl(t^{\frac{\ell}{d}}\bigr)$ for some $\ell\in\mathbb{Z}_{\geq 0}$. Then, we define $\operatorname{ord}_{\beta}(\mathfrak{a}):=\frac{\ell}{d}$.
@@
Since $\operatorname{Fitt}^{n}\Bigl(\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)=\operatorname{Jac}_{\overline{B}^{(\gamma)}/k[t]}\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}$, we have
@@
$\operatorname{ord}_{\beta}\left(\operatorname{Fitt}^{n}\Bigl(\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)\right)=\operatorname{ord}_{\alpha}\left(\operatorname{Jac}_{\overline{B}^{(\gamma)}/k[t]}\right)<\infty.$
Therefore, $\overline{B}^{(\gamma)}_{1/d}$ is smooth over $k[t^{1/d}]$ at the point $\beta(\eta)$, where $\eta$ denotes the generic point of $\operatorname{Spec}(k[[t^{1/d}]])$. Hence we have an isomorphism of $k[[t^{1/d}]]$-modules:
$\beta^{*}\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\big/\Bigl(\beta^{*}\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)_{\rm tor}\simeq k[[t^{1/d}]]^{\oplus n}.$
Let $a$ be the length of the $k[[t^{1/d}]]$-module
@@
$\operatorname{Coker}\left(\beta^{*}\Bigl(\Omega_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\to\beta^{*}\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\big/\Bigl(\beta^{*}\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)_{\rm tor}\right).$
Taking the $n$-th exterior power, the isomorphism (i) implies that $a$ is equal to the length of
@@
$\operatorname{Coker}\left(\beta^{*}\Bigl(\Omega^{n}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\to\beta^{*}\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\big/\Bigl(\beta^{*}\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)_{\rm tor}\right).$
We shall prove
$\frac{a}{d}+\operatorname{ord}_{\beta}\left(\operatorname{Fitt}^{n}\Bigl(\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)\right)=\operatorname{ord}_{\beta}\Bigl(\mathfrak{n}^{\prime}_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)$.
By the definition of $\mathfrak{n}^{\prime}_{\overline{B}^{(\gamma)}_{1/d}}$, we have
$\operatorname{ord}_{\beta}\Bigl(\mathfrak{n}^{\prime}_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)=\frac{1}{d}\operatorname{length}\left(\operatorname{Coker}\Bigl(\beta^{*}\Bigl(\Omega^{n}_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\to\beta^{*}L_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\right)$.
By Lemma 4.3(1), we have
$\operatorname{ord}_{\beta}\left(\operatorname{Fitt}^{n}\Bigl(\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)\right)=\frac{1}{d}\operatorname{length}\left(\operatorname{Coker}\Bigl(\beta^{*}\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\to\beta^{*}L_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\right)$.
Since $L_{\overline{B}^{(\gamma)}_{1/d}}$ is torsion-free, the map $\beta^{*}\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\to\beta^{*}L_{\overline{B}^{(\gamma)}_{1/d}}$ factors through the torsion-free quotient
$\beta^{*}\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\big/\Bigl(\beta^{*}\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)_{\rm tor}.$
Therefore, $\operatorname{ord}_{\beta}\left(\operatorname{Fitt}^{n}\Bigl(\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)\right)$ is equal to
$\frac{1}{d}\operatorname{length}\left(\operatorname{Coker}\Bigl(\beta^{*}\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\big/\Bigl(\beta^{*}\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)_{\rm tor}\to\beta^{*}L_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\right)$.
Thus, the assertion (iii) follows from (ii), (iv), and (v).
Since
$\operatorname{Coker}\left(\beta^{*}\Bigl(\Omega_{B_{k[t^{1/d}]}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\Bigr)\to\Bigl(\beta^{*}\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)\big/\Bigl(\beta^{*}\Omega_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\Bigr)_{\rm tor}\right)$
@@
we have $\frac{a}{d}=\operatorname{ord}_{\alpha}\bigl(\operatorname{jac}_{\overline{\mu}_{\gamma}}\bigr)$ . Therefore, the assertion follows from (iii) and Lemma 4.5. ∎
We define a $k[t]$-scheme $\widetilde{B}^{(\gamma)}_{1/d}$ by
$\widetilde{B}^{(\gamma)}_{1/d}$ $:=\operatorname{Spec}\left(R^{C_{\gamma}}[t^{1/d}]\big/\left(I_{\overline{B}^{(\gamma)}}\cap R^{C_{\gamma}}[t]\right)R^{C_{\gamma}}[t^{1/d}]\right)$
@@
Then, we have a natural morphism $\overline{B}^{(\gamma)}_{1/d}\to\widetilde{B}^{(\gamma)}_{1/d}$.
@@
By composing $\Omega^{n}_{\widetilde{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}$ and $\Omega^{n}_{\overline{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\to L_{\overline{B}^{(\gamma)}_{1/d}}$ in Lemma 4.3(1), we obtain a natural morphism $\Omega^{n}_{\widetilde{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to L_{\overline{B}^{(\gamma)}_{1/d}}$. Then, we denote by $\mathfrak{n}^{\prime\prime}_{\overline{B}^{(\gamma)}_{1/d}}$ the ideal sheaf on $\overline{B}^{(\gamma)}_{1/d}$ satisfying
$\operatorname{Im}\left(\Omega^{n}_{\widetilde{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to L_{\overline{B}^{(\gamma)}_{1/d}}\right)=\mathfrak{n}^{\prime\prime}_{\overline{B}^{(\gamma)}_{1/d}}\otimes_{\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}}L_{\overline{B}^{(\gamma)}_{1/d}}$.
@@
There exists a $C_{\gamma}$-invariant ideal sheaf $\mathfrak{n}^{\prime}_{p}$ on $\overline{B}^{(\gamma)}$ such that $\mathfrak{n}^{\prime\prime}_{\overline{B}^{(\gamma)}_{1/d}}=\mathfrak{n}^{\prime}_{p}\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}$.
@@
Take elements $g_{1},\ldots,g_{\ell}\in R^{C_{\gamma}}$ satisfying $R^{C_{\gamma}}=k[g_{1},\ldots,g_{\ell}]$. For a subset $J\subset\{1,\ldots,\ell\}$ with #J=Nc\#J=N-c, we define $\Delta_{J}\in R[t]$ as the determinant of the Jacobian matrix (of size $N$) with respect to $\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}$ for $1\leq i\leq c$ and $g_{i}$ for $i\in J$, and $\partial x_{j}$ for $1\leq j\leq N$. Note that $\Delta_{J}$ is defined up to sign. We define an ideal $\mathfrak{n}^{\prime}_{p}\subset R[t]/I_{\overline{B}^{(\gamma)}}$ as
@@
$\mathfrak{n}^{\prime}_{p}:=\bigl(\overline{\Delta_{J}}\ \big|\ \text{$J\subset\{1,\ldots,\ell\}$ with $\#J=N-c$}\bigr),$
where $\overline{\Delta_{J}}$ denoted the image of $\Delta_{J}\in R[t]$ in $R[t]/I_{\overline{B}^{(\gamma)}}$. For any $\delta\in C_{\gamma}$, by Lemma 3.5(1), we have
$\delta\left(\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\right)=\xi^{dw_{\delta}(f_{i})}\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}.$
@@
Furthermore, note that $\delta$ maps $x_{i}$’s to their $k$-linear independent $k$-linear combinations. Therefore, for each $J$, we have
$\xi^{dw_{\delta}({\bf f})}\cdot\Delta_{J}\cdot dx_{1}\wedge\cdots\wedge dx_{N}$
@@
$=\xi^{dw_{\delta}({\bf f})}\cdot\bigl(\wedge_{i\in J}\ dg_{i}\bigr)\wedge\left(\wedge_{1\leq i\leq c}\ d\left(\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\right)\right)$
@@
$=\bigl(\wedge_{i\in J}\ d(\delta(g_{i}))\bigr)\wedge\left(\wedge_{1\leq i\leq c}\ d\left(\delta\left(\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\right)\right)\right)$
@@
$=\delta\left(\bigl(\wedge_{i\in J}\ dg_{i}\bigr)\wedge\left(\wedge_{1\leq i\leq c}\ d\left(\lambda^{*}_{\gamma}(f_{i})t^{-w_{\gamma}(f_{i})}\right)\right)\right)$
$=\delta\left(\Delta_{J}\cdot dx_{1}\wedge\cdots\wedge dx_{N}\right)$
$=\delta\left(\Delta_{J}\right)\cdot d(\delta(x_{1}))\wedge\cdots\wedge d(\delta(x_{N}))$
$=\delta\left(\Delta_{J}\right)\cdot s\cdot dx_{1}\wedge\cdots\wedge dx_{N}$
@@
for some $s\in k^{\times}$. Hence, we have $\delta(\Delta_{J})=s^{\prime}\cdot\Delta_{J}$ for some $s^{\prime}\in k$. Hence, we conclude that $\mathfrak{n}^{\prime}_{p}$ is $C_{\gamma}$-invariant.
@@
We shall prove that this $\mathfrak{n}^{\prime}_{p}$ satisfies the assertion $\mathfrak{n}^{\prime\prime}_{\overline{B}^{(\gamma)}_{1/d}}=\mathfrak{n}^{\prime}_{p}\mathcal{O}_{\overline{B}^{(\gamma)}_{1/d}}$. We set
$S_{\widetilde{B}^{(\gamma)}_{1/d}}:=R^{C_{\gamma}}[t^{1/d}]\big/\left(I_{\overline{B}^{(\gamma)}}\cap R^{C_{\gamma}}[t]\right)R^{C_{\gamma}}[t^{1/d}].$
@@
We also use the notation in the proof of Lemma 4.3. By the proof of Lemma 4.3(1), the considered map $\Omega^{n}_{\widetilde{B}^{(\gamma)}_{1/d}/k[t^{1/d}]}\big|_{\overline{B}^{(\gamma)}_{1/d}}\to L_{\overline{B}^{(\gamma)}_{1/d}}$ is corresponding to